OpenStudy (anonymous):

(x-5)^2=20 equations by sqaure root

OpenStudy (anonymous):

Take square root both sides, x-5 = 2squareroot5 x= 5 + 2square root5

OpenStudy (nenadmatematika):

x^2-10x+25-20=0 so x^2-10x+5=0 and solve for x